# On the nullspace of split graphs

### v1

@@ 4 Unbalanced split graphs
    If $x\in\operatorname{\mathcal{N}}(Sp)$, then $x_{u}=x_{v}$ for all $u,v\in W(Sp)$, by Lemma 4.5 and Lemma 4.4. Since $\operatorname{Supp}(Sp)$ is independent by Theorem 4.2, it must be $x_{w}=0$ for all $w\in W(Sp)$, because $|W(Sp)|\geq 2$ and $W(Sp)$ is a clique. Consequently, $W(Sp)\cap\operatorname{Supp}(Sp)=\varnothing$.
    Now, observe that $S\setminus W(Sp)$ is the set of vertices that always lie in the independent part of every s-partition. By Theorem 4.2, for any s-partition $(K^{\prime},S^{\prime})$ of $Sp$, we have $\operatorname{Supp}(Sp)\subset S^{\prime}$. In particular, for each $w\in W(Sp)$, there exists an s-partition where $S^{\prime}=(S\setminus W(Sp))\cup\{w\}$. Therefore,
    $\operatorname{Supp}(Sp)\subset\bigcap_{w\in W(Sp)}\left((S\setminus W(Sp))\cup\{w\}\right)=S\setminus W(Sp).$
@@
    If $u,v$ are distinct swing vertices, observe that $A(Sp)_{*u}=A(Sp)_{*v}$ (since they are twins and both in the independent part in some s-partition). As a consequence, $e_{u}-e_{v}\in\operatorname{\mathcal{N}}(Sp)$ ($e_{i}$ is the $i$-th canonical vector). Then, $W(Sp)\subset\operatorname{Supp}(Sp)$.
    On the other hand, $Sp$ has a unique maximum independent set $S_{\max}$ which must include $W(Sp)$ (since swing vertices can be in the independent part). Hence, $S_{\max}=S\cup W(Sp)$. By Theorem 4.2, $\operatorname{Supp}(Sp)\subset S_{\max}=S\cup W(Sp)$.
@@
Since $Sp$ is connected and has no isolated vertices in $S$, all the columns of $R$ are nonzero. Since $K=\bigcup_{v\in S}N_{Sp}(v)$, all the rows of $R$ are nonzero. Moreover, since $\operatorname{rk}(R)=1$ and $R_{ij}\in\{0,1\}$ for all $i,j$, all columns of $R$ must be scalar multiples of each other. Since all columns are nonzero $\{0,1\}$-vectors, they must all be equal to the same vector. Since $R_{i*}\neq 0$ for all $i$ and $R$ is a $\{0,1\}$-matrix, we get that $R_{*j}=\operatorname{1\!\!1}$ for all $j$. Therefore, $R=J$.
For the twin property: Since $R=J$, every vertex in $S$ is adjacent to all vertices in $K$, so all vertices in $S$ have the same neighborhood $N_{Sp}(v)=K$ for all $v\in S$, making them twins. Similarly, every vertex in $K$ is adjacent to all vertices in $S$ and all other vertices in $K$ (since $K$ is a clique), so all vertices in $K$ have the same neighborhood $N_{Sp}(u)=(K\setminus\{u\})\cup S$ for all $u\in K$, making them twins as well. ∎
@@ 5 Split graphs with supported clique vertices
This section delves into the detailed structure of the nullspace when it involves vertices from the clique, a scenario that introduces rich and constrained linear dependencies. We prove that the subspace of kernel vectors with support in the clique—the clique-kernel —is at most one-dimensional. This foundational result leads to a precise dichotomy: the overall nullity of the split graph is either equal to the nullity of its bipartite submatrix $R$, or exactly one greater. This establishes a complete and computable classification of the nullity based purely on the dimension of this key subspace, tightly linking the combinatorial presence of supported clique vertices to the algebraic property of the graph’s nullity.
@@
Let
$A(Sp)=\begin{pmatrix}J-I&R\\ R^{t}&0\end{pmatrix}$
be the adjacency matrix of a split graph $Sp$ with a bipartition $(K,S)$ such that $|K|\geq 2$. If
$\operatorname{Cker}{\left(Sp\right)}=\operatorname{\mathcal{N}}(R^{t})\cap\operatorname{\mathcal{R}}((I-J)^{-1}R)\neq\{0\},$
then
$\dim(\operatorname{Cker}{\left(Sp\right)})=1.$
@@
Since $\operatorname{Cker}{\left(Sp\right)}\neq 0$, it is clear that $R\neq 0$. Let $\{b_{1},\ldots,b_{r}\}$ be a basis for the subspace $\operatorname{Cker}{\left(Sp\right)}$. For each $i\in[r]$, we have
@@
Let
$A(Sp)=\begin{pmatrix}J-I&R\\ R^{t}&0\end{pmatrix}$
be the adjacency matrix of a split graph $Sp$ with a bipartition $(K,S)$ such that $|K|\geq 2$. If
$\operatorname{Cker}{\left(Sp\right)}=\operatorname{\mathcal{N}}(R^{t})\cap\operatorname{\mathcal{R}}((I-J)^{-1}R)\neq\{0\},$
then
$\operatorname{Supp}(Sp)\cap K\neq\varnothing$
if and only if
$\dim(\operatorname{Cker}{\left(Sp\right)})=1.$
@@
Since
$\operatorname{\mathcal{N}}(Sp)\supseteq\mathcal{R}=\{(0,x)^{t}:x\in\operatorname{\mathcal{N}}(R)\}$
by item (2) of Theorem 3.2, we can extend a basis $B_{\mathcal{R}}$ for $\mathcal{R}$ to a basis for $\operatorname{\mathcal{N}}(Sp)$ to generate those kernel vectors $(x_{K},x_{S})^{t}\notin\mathcal{R}$ with $x_{K}\neq 0$.
Let $\operatorname{Cker}{\left(Sp\right)}=\operatorname{\mathcal{N}}(R^{t})\cap\operatorname{\mathcal{R}}((I-J)^{-1}R)$. We know from Theorem 5.2 that $\operatorname{Cker}{\left(Sp\right)}=\langle z_{0}\rangle$ is one-dimensional, where $z_{0}=(I-J)^{-1}Ry_{0}$ for some $y_{0}\in\mathbb{R}^{|S|}$ with $y_{0}\notin\operatorname{\mathcal{N}}(R)$ (since $\operatorname{\mathcal{N}}((I-J)^{-1}R)=\operatorname{\mathcal{N}}(R)$ and $z_{0}\neq 0$).
By items (1) and (2) of Theorem 3.4, for each vector $s\in\operatorname{\mathcal{N}}(Sp)\setminus\mathcal{R}$, we have $s_{K}\in\operatorname{Cker}{\left(Sp\right)}$, so $s_{K}=cz_{0}$ for some $c\in\mathbb{R}\setminus\{0\}$. Moreover, there exists some $y\in\mathbb{R}^{|S|}$ such that $(z_{0},y)^{t}\in\operatorname{\mathcal{N}}(Sp)$.
Any such vector $s$ can be written as:
@@
for some $c\in\mathbb{R}\setminus\{0\}$ and $x_{0}\in\operatorname{\mathcal{N}}(R)$, where we fix a particular representative $y_{0}$ such that $(z_{0},y_{0})^{t}\in\operatorname{\mathcal{N}}(Sp)$.
Therefore, $B_{\mathcal{R}}\dot{\cup}\{(z_{0},y_{0})^{t}\}$ is a basis for $\operatorname{\mathcal{N}}(Sp)$, and since $|B_{\mathcal{R}}|=\operatorname{nl}(R)$, we conclude that $\operatorname{nl}(Sp)=\operatorname{nl}(R)+1$. ∎
@@
Let $Sp(K,S)$ be a split graph with $|K|\geq 2$ and let $\operatorname{Cker}{\left(Sp\right)}=\operatorname{\mathcal{N}}(R^{t})\cap\operatorname{\mathcal{R}}((I-J)^{-1}R)$. Then, $\dim(\operatorname{Cker}{\left(Sp\right)})\leq 1$ and
@@
By Lemma 5.1, $\dim(\operatorname{Cker}{\left(Sp\right)})\leq 1$.
If $\dim(\operatorname{Cker}{\left(Sp\right)})=0$, then by Theorem 3.4(3), $\operatorname{Supp}(Sp)\subset S$, and hence $\operatorname{nl}(Sp)=\operatorname{nl}(R)$ by Theorem 3.2(3).
If $\dim(\operatorname{Cker}{\left(Sp\right)})=1$, then by Theorem 5.2, $\operatorname{Supp}(Sp)\cap K\neq\varnothing$, and hence $\operatorname{nl}(Sp)=\operatorname{nl}(R)+1$ by Theorem 5.3. ∎
@@ 6 Split graphs with nullity 1
Having established the fundamental structure of the nullspace for general split graphs, we now focus on the case where the nullity is exactly one. This condition imposes a remarkable rigidity on the graph’s structure, forcing a highly constrained relationship between the kernel, the adjugate of the adjacency matrix, and the graph’s support. In this section, we leverage the classical properties of the adjugate for singular matrices to derive precise combinatorial descriptions of these graphs, culminating in formulas that link the coordinates of the unique kernel vector to the determinants of principal submatrices.
Figure 3: Split graph $S$ with supported clique vertices and $\operatorname{nl}(S)=1$.
@@
Figure 4: Representation of the split graph $S$ without clique edges. The upper row corresponds to II and the lower row to $K$ (black vertices). Since $\operatorname{\mathcal{N}}(S)=\langle z\rangle$, the coordinates of $z$ are shown.
@@
Figure 5: Bipartite representation of a split graph $S$ (clique vertices in black) with $\operatorname{nl}(S)=1$. The numbers next to each vertex indicate the coordinates of “the” spanning kernel vector.
Lemma 6.1 ([horn2012matrix]).
@@
While this is a standard result (see, [horn2012matrix]), we include a concise proof here as this lemma plays a crucial role in our characterization of split graphs with nullity one, and we wish to keep our presentation self-contained.
@@
Let $n$ be the order of $A$. Recall that $\operatorname{rk}(\operatorname{adj}(A))$ is $n$ if $\operatorname{rk}(A)=n$, $1$ if $\operatorname{rk}(A)=n-1$, and 0 if $\operatorname{rk}(A)\leq n-2$. Thus $\operatorname{adj}(A)\neq 0$ if and only if $\operatorname{rk}(A)\geq n-1$, which is equivalent to $\operatorname{nl}(A)\leq 1$. ∎
@@
Let $n$ be the order of $A$ and let $\operatorname{adj}(A)=[a_{1}\mid\ldots\mid a_{n}]$, where $a_{i}$ denotes the $i$-th column.
Since $\operatorname{nl}(A)=1$, we have $\det(A)=0$ and $\operatorname{rk}(A)=n-1$. The fundamental identity $A\operatorname{adj}(A)=\det(A)I=0$ implies that
$A\operatorname{adj}(A)=[Aa_{1}\mid\ldots\mid Aa_{n}]=0,$
so $a_{i}\in\operatorname{\mathcal{N}}(A)$ for all $i\in[n]$.
Since $\operatorname{rk}(A)=n-1$, there exists an invertible $(n-1)\times(n-1)$ submatrix $A^{\prime}$ of $A$. The corresponding cofactor (which appears as an entry in $\operatorname{adj}(A)$) is $\pm\det(A^{\prime})\neq 0$, showing that $\operatorname{adj}(A)\neq 0$. Therefore, $\operatorname{rk}(\operatorname{adj}(A))=1$. ∎
If $A$ is a symmetric matrix, then $\operatorname{adj}(A)$ is symmetric as well.
@@
    If $(x,0)^{t}\in\operatorname{\mathcal{N}}(G)$, with $x\in\mathbb{R}^{s}\setminus\{0\}$, then
@@
    Let $\operatorname{\mathcal{N}}(G)=\langle z\rangle$ and $s=|\operatorname{Supp}(G)|$. First, we label $V(G)$ using the ordering $(\operatorname{Supp}(G),V(G)-\operatorname{Supp}(G))$. By this labeling, $z=(x,0)^{t}$, where $x\in\mathbb{R}^{s}$ and $x_{v}\neq 0$ for all $v\in[s]$. Then, by Lemma 6.2, we can write
@@
    Since $A(G)$ is symmetric, $\operatorname{adj}(G)$ is symmetric, so $Z$ must be symmetric. From the structure $Z=x(1,c_{2},\ldots,c_{s})$, symmetry implies that $c_{i}=\frac{x_{i}}{x_{1}}$ for all $i\in[s]$. Therefore, $Z=x_{1}^{-1}xx^{t}$.
@@
Let $Sp(K,S)$ be a split graph such that $\operatorname{Supp}(Sp)\cap K\neq\varnothing$, with $|K|\geq 2$. Then, $\operatorname{nl}(Sp)=1$ if and only if $\operatorname{\mathcal{N}}(R)=0$.
@@
Let $A$ be an $m\times n$ matrix such that $n\geq 2$, $\operatorname{nl}(A)=1$, and $A_{*j}\neq 0$ for all $j\in[n]$. If $z\in\operatorname{\mathcal{N}}(A)\setminus\{0\}$ and $z$ has no zero entries, then every set of n1n-1 columns of $A$ is a basis for $\operatorname{\mathcal{R}}(A)$.
@@
Since $z\in\operatorname{\mathcal{N}}(A)$ has no zero entries, for each $k\in[n]$ we have $A_{*k}=-\frac{1}{z_{k}}\sum_{j\neq k}z_{j}A_{*j}$, so each $B_{i}=\{A_{*j}:j\neq i\}$ spans $\operatorname{\mathcal{R}}(A)$. Since $\dim(\operatorname{\mathcal{R}}(A))=n-1$ and $|B_{i}|=n-1$, each $B_{i}$ is a basis. ∎
@@
Let $s=|\operatorname{Supp}(A)|\geq 2$. Consider the restriction of $A$ to the columns in $\operatorname{Supp}(A)$, and let $z_{\operatorname{Supp}}$ be the restriction of a nonzero kernel vector $z$ to these coordinates. Since $z_{j}\neq 0$ for all $j\in\operatorname{Supp}(A)$, we can apply Lemma 6.5 to conclude that $B_{i}$ spans $\langle\{A_{*j}:j\in\operatorname{Supp}(A)\}\rangle$.
The set $B$ contains: - $B_{i}$: s1s-1 vectors spanning the columns from $\operatorname{Supp}(A)$ - $\{A_{*j}:j\notin\operatorname{Supp}(A)\}$: nsn-s vectors
Thus $|B|=(s-1)+(n-s)=n-1=\operatorname{rk}(A)$. Since $B$ spans $\operatorname{\mathcal{R}}(A)$ and has the correct cardinality, it is a basis for $\operatorname{\mathcal{R}}(A)$. ∎
@@
is a basis for $\operatorname{\mathcal{R}}(A(G))$.
@@
Since $G$ has no isolated vertices, $A(G)_{*j}\neq 0$ for all $j\in V(G)$. The condition $\operatorname{nl}(G)=1$ implies $|\operatorname{Supp}(G)|\geq 2$ (otherwise the kernel would be trivial). Therefore, we can apply Lemma 6.6 with $A=A(G)$ and $i=s$. ∎
@@
Let $Sp(K,S)$ be a split graph such that
$\operatorname{Cker}{\left(Sp\right)}=\operatorname{\mathcal{N}}(R^{t})\cap\operatorname{\mathcal{R}}((I-J)^{-1}R)\neq 0,$
and let $P$ be the submatrix of $R$ of size $|K|\times|S^{\prime}|$ whose columns form a basis for $\operatorname{\mathcal{R}}(R)$. If $(Sp^{\prime},K,S^{\prime})$ is the split graph associated with $P$, then
$\operatorname{Cker}{\left(Sp\right)}=\operatorname{\mathcal{N}}(P^{t})\cap\operatorname{\mathcal{R}}((I-J)^{-1}P)$
@@
Since $P$ is obtained by selecting a basis for $\operatorname{\mathcal{R}}(R)$, we have: $\operatorname{\mathcal{R}}(P)=\operatorname{\mathcal{R}}(R)$, $\operatorname{\mathcal{N}}(P^{t})=\operatorname{\mathcal{N}}(R^{t})$, $\operatorname{\mathcal{N}}(P)=0$.
Therefore, $\operatorname{Cker}{\left(Sp\right)}=\operatorname{\mathcal{N}}(P^{t})\cap\operatorname{\mathcal{R}}((I-J)^{-1}P)$ remains unchanged.
Since $\operatorname{Cker}{\left(Sp\right)}\neq 0$, by Theorem 3.4(3), we have $\operatorname{Supp}(Sp^{\prime})\cap K\neq\varnothing$. Since $\operatorname{\mathcal{N}}(P)=0$, applying Corollary 6.4 to $Sp^{\prime}$ gives $\operatorname{nl}(Sp^{\prime})=1$. ∎
@@
Let $Sp(K,S)$ be a split graph such that
$\operatorname{Cker}{\left(Sp\right)}=\operatorname{\mathcal{N}}(R^{t})\cap\operatorname{\mathcal{R}}((I-J)^{-1}R)=\langle z\rangle,$
for some $z\in\mathbb{R}^{|K|}\setminus\{0\}$. Assume that the first $\operatorname{rk}(R)$ columns of $R$ are linearly independent and let $P$ be the submatrix of $R$ of size $|K|\times|S^{\prime}|$ obtained by deleting the last $|S|-\operatorname{rk}(R)$ columns of $R$. If $(Sp^{\prime},K,S^{\prime})$ is the induced split subgraph of $Sp$ associated with $P$, then we have the following:
@@
    $B_{s}=\{A(Sp)_{*j}:j\in V(Sp^{\prime})\setminus\{s\}\}$
    is a basis for $\operatorname{\mathcal{R}}(A(Sp))$.
@@
    $\operatorname{rk}(A(Sp))=|V(Sp^{\prime})|-1=\operatorname{rk}(R)+|K|-1$.
@@
    is a basis for $\operatorname{\mathcal{N}}(Sp)$ ($e_{v}$ is the $v$-th canonical vector).

### v2

@@ 4 Unbalanced split graphs
    From Theorem 4.2(2) we know that $\mathcal{S}(Sp)$ contains all the maximum independent sets of $Sp$. Thus, from Theorem 4.1(2) we easily derive that the intersection of all of them is $S^{*}(Sp)$. Hence, $\operatorname{Supp}(Sp)\subset S^{*}(Sp)$ by Theorem 4.4, and so $W(Sp)\cap\operatorname{Supp}(Sp)=\varnothing$ by (2).
    Alternatively: if $x\in\operatorname{\mathcal{N}}(Sp)$, then $x_{u}=x_{v}$ for all $u,v\in W(Sp)$, by Lemma 4.7 and Lemma 4.6. Since $\operatorname{Supp}(Sp)$ is independent by Theorem 4.4, it must be $x_{w}=0$ for all $w\in W(Sp)$, because $|W(Sp)|\geq 2$ and $W(Sp)$ is a clique. Consequently, $W(Sp)\cap\operatorname{Supp}(Sp)=\varnothing$.
@@
    If $u,v$ are distinct swing vertices, observe that $A(Sp)_{*u}=A(Sp)_{*v}$. As a consequence, $e_{u}-e_{v}\in\operatorname{\mathcal{N}}(Sp)$. Then, $W(Sp)\subset\operatorname{Supp}(Sp)$.
    On the other hand, we know from Theorem 4.1(3) and Theorem 4.2(2) that $S^{*}(Sp)\cup W(Sp)$ is the unique maximum independent set of $Sp$. Therefore, $\operatorname{Supp}(Sp)\subset S^{*}(Sp)\cup W(Sp)$, by Theorem 4.4.
@@
Since $Sp$ is connected and has no isolated vertices in $S$, all the columns of $R$ are nonzero. Since $K=\bigcup_{v\in S}N(v)$, all the rows of $R$ are nonzero. Moreover, since $\operatorname{rk}(R)=1$, all the columns of $R$ must be scalar multiples of each other. Since all columns are nonzero $\{0,1\}$-vectors, they must all be equal to the same vector. Since $R_{i*}\neq 0$ for all $i$ and $R$ is a $\{0,1\}$-matrix, we get that $R_{*j}=\operatorname{1\!\!1}$ for all $j$. Therefore, $R=J$.
Since $R=J$, every vertex in $S$ is adjacent to all vertices in $K$, i.e., $N(v)=K$ for all $v\in S$. Therefore, all the independent vertices are pairwise twins. Similarly, every vertex in $K$ is adjacent to all vertices in $S$ and all other vertices in $K$. Hence,
$N(u)\setminus\{v\}=(K\setminus\{u,v\})\cup S=N(v)\setminus\{u\}$
for all $u,v\in K$, which means that all the clique vertices are pairwise twins. ∎
@@ 5 Split graphs with supported clique vertices
In this section we examine the structure of the nullspace when clique vertices are involved; this scenario gives rise to rich yet highly constrained linear dependencies. We prove that the subspace of kernel vectors with support in the clique—the clique-kernel —is at most one-dimensional. This foundational result leads to a precise dichotomy: the overall nullity of the split graph is either equal to the nullity of its biadjacency matrix $R$, or exactly one greater. This establishes a complete and computable classification of the nullity based purely on the dimension of this key subspace, tightly linking the combinatorial presence of supported clique vertices to the algebraic property of the graph’s nullity.
@@
Let $Sp(K,S)$ be a split graph such that $|K|\geq 2$. If $\operatorname{Cker}{\left(Sp\right)}\neq\{0\}$, then $\dim(\operatorname{Cker}{\left(Sp\right)})=1$
@@
Since $\operatorname{Cker}{\left(Sp\right)}\neq\{0\}$, it is clear that $R\neq 0$. Let $\{b_{1},\ldots,b_{r}\}$ be a basis for the subspace $\operatorname{Cker}{\left(Sp\right)}$. For each $i\in[r]$, we have
@@
Let $Sp(K,S)$ be a split graph such that $|K|\geq 2$. If $\operatorname{Cker}{\left(Sp\right)}\neq\{0\}$, then $\operatorname{Supp}(Sp)\cap K\neq\varnothing$ if and only if $\dim(\operatorname{Cker}{\left(Sp\right)})=1$.
@@
Let $\mathcal{X}=\{(0,x)^{t}:x\in\operatorname{\mathcal{N}}(R)\}$. Since $\mathcal{X}\subset\operatorname{\mathcal{N}}(Sp)$ by Theorem 3.2(2), we can extend a basis $B_{\mathcal{X}}$ for $\mathcal{X}$ to a basis for $\operatorname{\mathcal{N}}(Sp)$ to span those kernel vectors $(x_{K},x_{S})^{t}$ with $x_{K}\neq 0$.
We know from Theorem 5.2 that $\operatorname{Cker}{\left(Sp\right)}=\langle z_{0}\rangle$, where $z_{0}=(I-J)^{-1}Ry$, for some $y$. Since $\operatorname{\mathcal{N}}((I-J)^{-1}R)=\operatorname{\mathcal{N}}(R)$ and $z_{0}\neq 0$, such a vector $y$ belongs to $y_{0}+\operatorname{\mathcal{N}}(R)$, where $y_{0}$ is fixed and $y_{0}\notin\operatorname{\mathcal{N}}(R)$. Then, by items (1) and (2) of Theorem 3.4, every vector $s\in\operatorname{\mathcal{N}}(Sp)\setminus\mathcal{X}$ has the form
@@
for some $c\in\mathbb{R}\setminus 0$ and some $x_{0}\in\operatorname{\mathcal{N}}(R)$. Therefore, $B_{\mathcal{X}}\dot{\cup}\{(z_{0},y_{0})^{t}\}$ is a basis for $\operatorname{\mathcal{N}}(Sp)$. ∎
@@
If $Sp(K,S)$ is a split graph with $|K|\geq 2$, then $\dim(\operatorname{Cker}{\left(Sp\right)})$ $\leq 1$ and
@@
By Lemma 5.1, $\dim(\operatorname{Cker}{\left(Sp\right)})\leq 1$. If $\dim(\operatorname{Cker}{\left(Sp\right)})=0$, then by Theorem 3.4(3), $\operatorname{Supp}(Sp)\subset S$, and hence $\operatorname{nl}(Sp)=\operatorname{nl}(R)$ by Theorem 3.2(3). If $\dim(\operatorname{Cker}{\left(Sp\right)})=1$, then by Theorem 5.2, $\operatorname{Supp}(Sp)\cap K\neq\varnothing$, and hence $\operatorname{nl}(Sp)=\operatorname{nl}(R)+1$ by Theorem 5.3. ∎
@@ 6 Split graphs with nullity 1
Having determined the general structure of the nullspace, we now specialize in split graphs whose nullity is exactly one. This condition imposes a remarkable rigidity on the graph’s structure, forcing a highly constrained relationship between the kernel, the adjugate of the adjacency matrix, and the graph’s support. In this section, we leverage the classical properties of the adjugate for singular matrices to derive precise combinatorial descriptions of these graphs, culminating in formulas that link the coordinates of the unique kernel vector to the determinants of principal submatrices.
Figure 3: Split graph with supported clique vertices and nullity one.
@@
Figure 4: Bipartite representation of the split graph $Sp(K,S)$ of Figure 3. The upper row corresponds to $S$ and the lower row to $K$ (black vertices). Since $\operatorname{\mathcal{N}}(Sp)=\langle z\rangle$, the coordinates of $z$ are shown.
@@
Figure 5: Bipartite representation of a split graph (clique vertices in black) with nullity one. The numbers next to each vertex indicate the coordinates of “the” spanning kernel vector.
Lemma 6.1 ([10]).
@@
While this is a standard result (see [10]), we include a concise proof here as this lemma plays a crucial role in our characterization of split graphs with nullity one, and we wish to keep our presentation self-contained.
@@
Let $n$ be the order of $A$. The fact that $\operatorname{adj}(A)$ has at least a nonzero entry means that $A$ contains a nonzero minor of order n1n-1. This is equivalent to $\operatorname{rk}(A)\geq n-1$, because $\operatorname{rk}(A)$ is, by definition, the maximum size of an invertible submatrix of $A$. ∎
@@
Let $n$ be the order of $A$ and let $\operatorname{adj}(A)=[a_{1}\mid\ldots\mid a_{n}]$, where $a_{i}$ denotes the $i$-th column. Since $\det(A)=0$, the fundamental identity $\det(A)I=A\operatorname{adj}(A)$ implies that
$0=A\operatorname{adj}(A)=[Aa_{1}\mid\ldots\mid Aa_{n}].$
So, $a_{i}\in\operatorname{\mathcal{N}}(A)$ for all $i\in[n]$, from which we deduce that $\operatorname{rk}(\operatorname{adj}(A))\leq 1$, because $\operatorname{nl}(A)=1$. On the other hand, we know by Lemma 6.1 that $\operatorname{rk}(\operatorname{adj}(A))\neq 0$. Therefore, $\operatorname{rk}(\operatorname{adj}(A))=1$. ∎
If $A$ is a symmetric matrix, recall that $\operatorname{adj}(A)$ is symmetric as well.
@@
    If $(x,0)^{t}\in\operatorname{\mathcal{N}}(G)$, with $x\in\mathbb{R}^{s}\setminus 0$, then
@@
    Let $\operatorname{\mathcal{N}}(G)=\langle z\rangle$ and $s=|\operatorname{Supp}(G)|$. First, we label $V(G)$ using the ordering $(\operatorname{Supp}(G),V(G)\setminus\operatorname{Supp}(G))$. By this labeling, $z=(x,0)^{t}$, where $x\in\mathbb{R}^{s}$ and $x_{v}\neq 0$ for all $v\in[s]$. Then, by Lemma 6.2, we can write
@@
    Since $A(G)$ is symmetric, $\operatorname{adj}(G)$ is symmetric, so $Z$ must be symmetric. From the structure of $Z=x(1,c_{2},\ldots,c_{s})$, symmetry implies that $c_{i}=\frac{x_{i}}{x_{1}}$ for all $i\in[s]$. Therefore, $Z=x_{1}^{-1}xx^{t}$.
@@
Let $Sp(K,S)$ be a split graph such that $\operatorname{Supp}(Sp)\cap K\neq\varnothing$, with $|K|\geq 2$. Then, $\operatorname{nl}(Sp)=1$ if and only if $\operatorname{\mathcal{N}}(R)=\{0\}$.
@@
Let $A$ be an $m\times n$ matrix such that $n\geq 2$, $\operatorname{nl}(A)=1$, and $A_{*j}\neq 0$ for all $j\in[n]$. If $z\in\operatorname{\mathcal{N}}(A)\setminus 0$ and $z$ has no zero entries, then every set of n1n-1 columns of $A$ is a basis for $\operatorname{\mathcal{R}}(A)$.
@@
Since $z\in\operatorname{\mathcal{N}}(A)$ has no zero entries, for each $k\in[n]$ we have $A_{*k}=-\frac{1}{z_{k}}\sum_{j\neq k}z_{j}A_{*j}$, so each $B_{i}=\{A_{*j}:j\neq i\}$ spans $\operatorname{\mathcal{R}}(A)$. Since $\operatorname{rk}(A)=n-1=|B_{i}|$, each $B_{i}$ is a basis for $\operatorname{\mathcal{R}}(A)$. ∎
@@
From Lemma 6.5, we know that $B_{i}$ spans the subspace $\langle\{A_{*j}:j\in\operatorname{Supp}(A)\}\rangle$. With this in mind, it is clear that $B$ spans $\operatorname{\mathcal{R}}(A)$. Since $|B|=n-1=\operatorname{rk}(A)$, we can finally conclude that $B$ is a basis for $\operatorname{\mathcal{R}}(A)$. ∎
@@
is a basis for $\operatorname{\mathcal{R}}(G)$.
@@
Since $G$ has no isolated vertices, $A(G)_{*j}\neq 0$ for all $j\in V(G)$. If $\operatorname{Supp}(G)=\{u\}$ for some $u$, then $\operatorname{\mathcal{N}}(G)=\langle e_{u}\rangle$. But, $0=A(G)e_{u}=A(G)_{*u}$, contradicting that all the columns of $A(G)$ are nonzero. Therefore, $|\operatorname{Supp}(G)|\geq 2$, and so we can apply Lemma 6.6. ∎
@@
Let $Sp(K,S)$ be a split graph such that $\operatorname{Cker}{\left(Sp\right)}\neq\{0\}$, and let $P$ be the submatrix of $R$ of size $|K|\times|\operatorname{rk}(R)|$ whose columns form a basis for $\operatorname{\mathcal{R}}(R)$. Let $S^{\prime}$ be the set of vertices associated with the columns of $P$. If $Sp^{\prime}(K,S^{\prime})$ is the split graph whose biadjacency matrix is $P$, then
$\operatorname{Cker}{\left(Sp\right)}=\operatorname{Cker}{\left(Sp^{\prime}\right)}$
@@
Since $P$ is obtained by selecting a basis for $\operatorname{\mathcal{R}}(R)$, we have $\operatorname{\mathcal{R}}(P)=\operatorname{\mathcal{R}}(R)$ (hence, $\operatorname{\mathcal{R}}((I-J)^{-1}P)=\operatorname{\mathcal{R}}((I-J)^{-1}R)$) and $\operatorname{\mathcal{N}}(P^{t})=\operatorname{\mathcal{N}}(R^{t})$. Therefore, $\operatorname{Cker}{\left(Sp\right)}=\operatorname{Cker}{\left(Sp^{\prime}\right)}$. Since $\operatorname{Cker}{\left(Sp\right)}\neq\{0\}$, by Theorem 3.4(3), we have $\operatorname{Supp}(Sp^{\prime})$ $\cap K\neq\varnothing$. Since $\operatorname{\mathcal{N}}(P)=\{0\}$, applying Corollary 6.4 to $Sp^{\prime}$ gives $\operatorname{nl}(Sp^{\prime})=1$. ∎
@@
Let $Sp(K,S)$ be a split graph such that $\operatorname{Cker}{\left(Sp\right)}=\langle z\rangle$, for some $z\in\mathbb{R}^{|K|}\setminus 0$. Assume that the first $\operatorname{rk}(R)$ columns of $R$ are linearly independent and let $P$ be the submatrix of $R$ obtained by deleting the last $|S|-\operatorname{rk}(R)$ columns of $R$. Let $S^{\prime}$ be the set of vertices associated with the columns of $P$. If $Sp^{\prime}(K,S^{\prime})$ is the induced split subgraph of $Sp$ whose biadjacency matrix is $P$, then we have the following:
@@
    $B_{s}=\{A(Sp)_{*j}:j\in V(Sp^{\prime})\setminus s\}$
    is a basis for $\operatorname{\mathcal{R}}(Sp)$.
@@
    $\operatorname{rk}(Sp)=|V(Sp^{\prime})|-1=\operatorname{rk}(R)+|K|-1$.
@@
    is a basis for $\operatorname{\mathcal{N}}(Sp)$.